OpenStudy (anonymous):

m=-7(-2,-8) Write the equation on the form Ax+By=C

OpenStudy (anonymous):

y+8=-7(x+2)

OpenStudy (anonymous):

(y-(-8)) = -7(x-(-2)) Fool around a bit and you get y+7x=-22